Shift report 20011023 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Oct 23 2001 - 01:16:39 EDT

  • Next message: Apache: "Shift report 20011023 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Oct 23 01:16:39 2001, Shift: 16:00-24:00
    Supervisor: DO, with: CEJ for the food!
    Log book#10 pages 50 - 51
    
    Shift Summary:
    ------------------------------------------------------------
    1.Machine status:
      Not much good to say: there was a store that started before
      I came with low ZDC rates (about 120 Hz). I asked MCR
      to steer for us but they were a bit reluctant for some
      reasons unknown to me. A funny guy from Phenix (the exp. 
      NOT the city :) called me and we decided to vote for a 
      dump around 17:45. Yet, MCR let the beams  until 18:20
      (they have some big problems with the famous Quench Link
      Interlock). Since then, nothing. A couple of injections
      that failed quickly. They finally made a ramp for conducting 
      hysteresis study.
    
    2.Achievements:
      Run 5482 ended 16:05, MRS 4B - FS1/2A 52k tr1 - 15.7k tr6
    
      Run 5483 16:08-16:36, MRS 4B - FS1/2A 28k tr1 - 8.7 tr6  
          stopped it because MRS 4B goal achieved
    
      Run 5484 16:38-16:43, small run with wrong current in D5
          (600 Amps instead of 500)
    
      Run 5485 16:44-18:42, MRS 3B - FS1/2A 54k tr1 - 16k tr6
          stopped cause beam dumped by MCR    
    
      Detector pedestal run 5486
      TPC      pedestal run 5487
    
    3.No detector problems at all (since we were in a stand-by
      mode most the of the shift).
      Flemming changed something in the FEH around 16:00: he
      added a capacitive decoupling to retime the CDF with 
      the BB but as far as I know, it was not a success and
      gave a bigger offset.     
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Oct 23 2001 - 01:17:15 EDT